1. Oxygen can be made in the laboratory by reacting sodium peroxide and water:

2 Na2O2 (s) + 2 H2O (l) 🡪 4 NaOH (s) + O2 (g)

ΔH° = -109.0 kJ: ΔG° = -148.4 kJ at 25°C 

  1. Calculate ΔS°.  Is the sign reasonable?
  2. Calculate S° for Na2O2 (s).
  3. Calculate ΔHf° for Na2O2 (s). 
1
Expert's answer
2021-01-13T13:52:23-0500

Oxygen can be made in the laboratory by reacting sodium peroxide and water:

2 Na2O2 (s) + 2 H2O (l) 🡪 4 NaOH (s) + O2 (g)

ΔH° = -109.0 kJ: ΔG° = -148.4 kJ at 25°C 

ΔG° = ΔH° - TΔS°

= -148 * 1000 = -109 * 1000 + 298 * ΔS°


= -39000 = 298 * ΔS°

= ΔS°= 39000 / 298 = 131 j / k . mol